Whammo’s income statement is as follows:

Sales (10,000 units)

$90,000

Less variable costs

- 38,000

Contribution margin

$52,000

Less fixed costs

- 24,000

Net income

$ 28,000

If sales volume increases by 10%, profits will:
Answer #1

Profit S 33,200 Working Notes: Sales Volume = 10,000 Revised Sales Volume = 10,000+(10.000*10%) = 10,000+1,000 = 11,000 Selli